# Deep Learning for Protein-Ligand Docking: Are We There Yet?

### v3

@@ Appendix D Metrics
As originally proposed by Robin et al. [2023], in this study, we adopt the PLI-specific variant of lDDT, which calculates lDDT scores to compare predicted and ground-truth protein-ligand complex structures following optimal structural alignment of the predicted and ground-truth protein-ligand binding pockets.
@@ Appendix E Documentation for datasets
Below, we provide detailed documentation for each dataset included in our benchmark, summarised in Table 1. Each dataset is freely available for download from the benchmark’s accompanying Zenodo data record [Morehead et al., 2024] under a CC-BY 4.0 license. In lieu of being able to create associated metadata for each of our macromolecular datasets using an ML-focused library such as Croissant [Akhtar et al., 2024] (due to file type compatibility issues), instead, we report structured metadata for our preprocessed datasets using Zenodo’s web user interface [Morehead et al., 2024]. Note that, for all datasets, we authors bear all responsibility in case of any violation of rights regarding the usage of such datasets.
E.1 Astex Diverse Set - Single-Ligand Docking (Difficulty: Easy)
@@
  • Collection For this dataset, which was originally compiled by Hartshorn et al. [2007], we adopt the version further prepared by Buttenschoen et al. [2024].
  • Composition The dataset consists of 85 (80) single-ligand protein complexes (for which we could obtain high-accuracy predicted protein structures using AlphaFold/ESMFold).
  • Hosting Our preprocessed version of the dataset (https://doi.org/10.5281/zenodo.11477766) can be downloaded from the benchmark’s Zenodo data record at https://zenodo.org/records/11477766/files/astex_diverse_set.tar.gz.
  • Licensing We have released our preprocessed version of the dataset under a CC-BY 4.0 license. The original dataset is available under a CC-BY 4.0 license on Zenodo [Buttenschoen et al., 2023].
@@
E.2 PoseBusters Benchmark Set - Single-Ligand Docking (Difficulty: Intermediate)
@@
  • Collection For this dataset, we adopt the version introduced by Buttenschoen et al. [2024].
  • Composition The dataset consists of 308 (280) single-ligand protein complexes (for which we could obtain high-accuracy predicted protein structures using AlphaFold/ESMFold).
  • Hosting Our preprocessed version of the dataset (https://doi.org/10.5281/zenodo.11477766) can be downloaded from the benchmark’s Zenodo data record at https://zenodo.org/records/11477766/files/posebusters_benchmark_set.tar.gz.
  • Licensing We have released our preprocessed version of the dataset under a CC-BY 4.0 license. The original dataset is available under a CC-BY 4.0 license on Zenodo [Buttenschoen et al., 2023].
@@
E.3 DockGen Set - Single-Ligand Docking (Difficulty: Challenging)
The DockGen dataset was originally developed for docking individual ligands to target proteins in the context of novel protein binding pockets. As such, this dataset is useful for evaluating how well each baseline method can generalize to distinctly different binding pockets compared to those on which it may have been trained.
@@ E.3 DockGen Set - Single-Ligand Docking (Difficulty: Challenging)
  • Collection For this dataset, we adopt the version introduced by Corso et al. [2024a].
@@
  • Hosting Our preprocessed version of the dataset (https://doi.org/10.5281/zenodo.11477766) can be downloaded from the benchmark’s Zenodo data record at https://zenodo.org/records/11477766/files/dockgen_set.tar.gz.
  • Licensing We have released our preprocessed version of the dataset under a CC-BY 4.0 license. The original dataset is available under an MIT license on Zenodo [Corso et al., 2024b].
@@
E.4 CASP15 Set - Multi-Ligand Docking (Difficulty: Challenging)
@@
  • Motivation Multi-ligand evaluation datasets for molecular docking provide the rare opportunity to assess how well baseline methods can model intricate protein-ligand interactions while avoiding troublesome protein-ligand and ligand-ligand steric clashes. Additionally, more accurate modeling of multi-ligand complexes in future works may lead to improved techniques for computational enzyme design and regulation [Stärk et al., 2023].
  • Collection For this dataset, we manually collect each publicly and privately available CASP15 protein-bound ligand complex structure compatible with protein-ligand (e.g., non-RNA) benchmarking.
  • Composition The dataset consists of 102 (86) fragment ligands contained within 19 (15) separate (publicly available) protein complexes, of which 6 (2) and 13 (2) of these complexes are single and multi-ligand complexes, respectively.
  • Hosting Our preprocessed version of (the publicly available version of) this dataset (https://doi.org/10.5281/zenodo.11477766) can be downloaded from the benchmark’s Zenodo data record at https://zenodo.org/records/11477766/files/casp15_set.tar.gz.
  • Licensing We have released our preprocessed version of the (public) dataset under a CC-BY 4.0 license. The original (public) dataset is free for download via the RCSB PDB [Bank, 1971].
@@
Figure 7: Astex, PoseBusters, DockGen, and CASP15 dataset analysis of protein-ligand interactions.
@@ Appendix F Analysis of protein-ligand interactions
Inspired by a similar analysis presented in the PoseCheck benchmark [Harris et al., 2023], in this section, we study the frequency of different types of protein-ligand interactions such as Van der Waals contacts and hydrophobic interactions occurring natively within the Astex Diverse, PoseBusters Benchmark, DockGen, and CASP15 datasets, respectively. In particular, these measures allow us to better understand the diversity of interactions each baseline method within the PoseBench benchmark is tasked to model, within the context of each test dataset. Figure 7 displays the results of this analysis.
@@
Overall, we find that the Astex Diverse, PoseBusters Benchmark, and DockGen datasets contain similar types and frequencies of interactions, with the PoseBusters Benchmark and DockGen datasets containing slightly more hydrogen bond acceptors ($\sim$3 vs 1) and Van der Waals contacts ($\sim$13 vs 8) on average compared to the Astex Diverse dataset. However, we note a significant difference in interaction types and frequencies between the CASP15 dataset and the three other datasets. Specifically, we find it contains a significantly higher proportion of hydrogen bond acceptors and donors ($\sim$40), Van der Waals contact ($\sim$200), and hydrophobic interactions ($\sim$15) on average. Particularly interesting to note is the CASP15 dataset’s bimodal distribution of hydrophobic interactions, suggesting that the dataset contains two primary classes of interacting ligands giving rise to hydrophobic interactions. One possible explanation for this phenomenon is that the CASP targets, in contrast to the Astex Diverse, PoseBusters Benchmark, and DockGen targets, consist of a variety of both organic (e.g., drug-like molecules) and inorganic (e.g., metal) cofactors.
Appendix G Additional results
@@
G.1 DockGen results
DockGen dataset. The DockGen dataset [Corso et al., 2024a] contains 189 diverse single-ligand protein complexes, each representing a novel type of protein-ligand binding pocket. This dataset can be considered the most difficult single-ligand benchmark set since its protein binding sites are distinctly different from those found in the training datasets of most deep learning-based docking methods to date.
@@ G.1 DockGen results
Benchmark results. Figures 8 and 9 show that DiffDock-L, NeuralPLexer, and Ensemble (Con) provide the best pocket generalization capabilities compared to all other baseline methods. In particular, Ensemble (Con) seems to provide the most stable performance in this setting. Moreover, the results for DiffDock-L-Vina and P2Rank-Vina suggest that DiffDock-L generalizes to new binding pockets slightly better than P2Rank for conventional docking with AutoDock-Vina. Additionally, DiffDock-L’s results with protein-flexible relaxation applied (i.e., DiffDock-L (Relax-P)) demonstrate that fixed-protein relaxation (albeit unideal from a theoretical e.g., protein side chain perspective [Wankowicz et al., 2022]) yields less accuracy degradation to DiffDock-L’s original ligand conformations compared to side chain-flexible relaxation. However, we note that none of the baseline methods generate any PB-valid ligand conformations, suggesting that such "correct" poses are approximately accurate yet physically implausible in certain measurable ways.
@@
G.2 Expanded Astex & PoseBusters results
G.2.1 Pocket-only PoseBusters results
Figures 10 and 11 illustrate the impact of reducing the binding pocket search space of each baseline docking method by providing each method with alternative versions of the predicted PoseBusters Benchmark protein structures that have been cropped to contain only ligand-interacting ($<$ 4 Å heavy atom distance) protein residues and their (7) sequence-adjacent neighbors. Overall, we find that performing such pocket-level docking increases the docking success rates and favorably narrows the ligand RMSD distributions of DiffDock-L, NeuralPLexer, RoseTTAFold-All-Atom, and Ensemble (Con), whereas for all other baselines, performance is either maintained or degraded marginally. This finding highlights that methods such as DiffDock-L, NeuralPLexer, and RoseTTAFold-All-Atom are better at identifying the correct local structural conformations of each ligand compared to other baseline methods such as FABind and DynamicBind.
G.2.2 Astex & PoseBusters RMSD results

### v4

@@ Appendix D Metrics
As originally proposed by Robin et al. (2023), in this study, we adopt the PLI-specific variant of lDDT, which calculates lDDT scores to compare predicted and ground-truth protein-ligand complex structures following optimal structural alignment of the predicted and ground-truth protein-ligand binding pockets.
@@ Appendix E Documentation for datasets
Below, we provide detailed documentation for each dataset included in our benchmark, summarised in Table 1. Each dataset is freely available for download from the benchmark’s accompanying Zenodo data record (Morehead et al., 2024) under a CC-BY 4.0 license. In lieu of being able to create associated metadata for each of our macromolecular datasets using an ML-focused library such as Croissant (Akhtar et al., 2024) (due to file type compatibility issues), instead, we report structured metadata for our preprocessed datasets using Zenodo’s web user interface (Morehead et al., 2024). Note that, for all datasets, we authors bear all responsibility in case of any violation of rights regarding the usage of such datasets.
(a) RMSD of AlphaFold 3’s predictions.
(b) TM-score of AlphaFold 3’s predictions.
Figure 7: Accuracy of AlphaFold 3’s predicted protein structures for the Astex Diverse dataset.
(a) RMSD of AlphaFold 3’s predictions.
(b) TM-score of AlphaFold 3’s predictions.
Figure 8: Accuracy of AlphaFold 3’s predicted protein structures for the PoseBusters Benchmark dataset.
(a) RMSD of AlphaFold 3’s predictions.
(b) TM-score of AlphaFold 3’s predictions.
Figure 9: Accuracy of AlphaFold 3’s predicted protein structures for the DockGen dataset.
(a) RMSD of AlphaFold 3’s predictions.
(b) TM-score of AlphaFold 3’s predictions.
Figure 10: Accuracy of AlphaFold 3’s predicted protein structures for the CASP15 dataset.
E.1 Astex Diverse Set - Single-Ligand Docking
(Difficulty: Easy)
@@ E.1 Astex Diverse Set - Single-Ligand Docking
  • Collection For this dataset, which was originally compiled by Hartshorn et al. (2007), we adopt the version further prepared by Buttenschoen et al. (2024).
  • Composition The dataset consists of 80 single-ligand protein complexes for which we could obtain high-accuracy predicted protein structures using AlphaFold 3. The accuracy of the AlphaFold 3-predicted structures is measured in terms of their RMSD and TM-score (Zhang & Skolnick, 2004) compared to the corresponding ground-truth (i.e., experimental) protein structures and is visualized in Figure 7. Notably, 79% of the predicted structures have an RMSD below 4 Å and a TM-score above 0.7, indicating that most of the dataset’s proteins have a reasonably accurate predicted structure.
  • Hosting Our preprocessed version of the dataset (https://doi.org/10.5281/zenodo.13858866) can be downloaded from the benchmark’s Zenodo data record at https://zenodo.org/records/13858866/files/astex_diverse_set.tar.gz.
  • Licensing We have released our preprocessed version of the dataset under a CC-BY 4.0 license. The original dataset is available under a CC-BY 4.0 license on Zenodo (Buttenschoen et al., 2023). The pre-holo-aligned protein structures predicted by AlphaFold 3 for this dataset (available on Zenodo (Morehead et al., 2024)) must only be used in accordance with the Terms of Service provided by the AlphaFold Server.
@@
E.2 PoseBusters Benchmark Set - Single-Ligand Docking
(Difficulty: Intermediate)
@@ E.2 PoseBusters Benchmark Set - Single-Ligand Docking
  • Collection For this dataset, we adopt the version introduced by Buttenschoen et al. (2024).
  • Composition The dataset consists of 280 single-ligand protein complexes for which we could obtain high-accuracy predicted protein structures using AlphaFold 3. The accuracy of the AlphaFold 3-predicted structures is measured in terms of their RMSD and TM-score (Zhang & Skolnick, 2004) compared to the corresponding ground-truth (i.e., experimental) protein structures and is visualized in Figure 8. Notably, 70% of the predicted structures have an RMSD below 4 Å and a TM-score above 0.7, indicating that most of the dataset’s proteins have a reasonably accurate predicted structure.
  • Hosting Our preprocessed version of the dataset (https://doi.org/10.5281/zenodo.13858866) can be downloaded from the benchmark’s Zenodo data record at https://zenodo.org/records/13858866/files/posebusters_benchmark_set.tar.gz.
  • Licensing We have released our preprocessed version of the dataset under a CC-BY 4.0 license. The original dataset is available under a CC-BY 4.0 license on Zenodo (Buttenschoen et al., 2023). The pre-holo-aligned protein structures predicted by AlphaFold 3 for this dataset (available on Zenodo (Morehead et al., 2024)) must only be used in accordance with the Terms of Service provided by the AlphaFold Server.
@@
E.3 DockGen Set - Single-Ligand Docking
(Difficulty: Challenging)
The DockGen dataset was originally developed for docking individual ligands to target proteins in the context of novel protein binding pockets. As such, this dataset is useful for evaluating how well each baseline method can generalize to distinctly different binding pockets compared to those on which it commonly may have been trained.
@@ E.3 DockGen Set - Single-Ligand Docking
  • Collection For this dataset, we adopt the version introduced by Corso et al. (2024a).
@@
  • Hosting Our preprocessed version of the dataset (https://doi.org/10.5281/zenodo.13858866) can be downloaded from the benchmark’s Zenodo data record at https://zenodo.org/records/13858866/files/dockgen_set.tar.gz.
  • Licensing We have released our preprocessed version of the dataset under a CC-BY 4.0 license. The original dataset is available under an MIT license on Zenodo (Corso et al., 2024b). The pre-holo-aligned protein structures predicted by AlphaFold 3 for this dataset (available on Zenodo (Morehead et al., 2024)) must only be used in accordance with the Terms of Service provided by the AlphaFold Server.
@@
E.4 CASP15 Set - Multi-Ligand Docking
(Difficulty: Challenging)
@@ E.4 CASP15 Set - Multi-Ligand Docking
  • Motivation Multi-ligand evaluation datasets for molecular docking provide the rare opportunity to assess how well baseline methods can model intricate protein-ligand interactions while avoiding troublesome protein-ligand and ligand-ligand steric clashes. Additionally, more accurate modeling of multi-ligand complexes in future works may lead to improved techniques for computational enzyme design and regulation (Stärk et al., 2023).
  • Collection For this dataset, we manually collect each publicly and privately available CASP15 protein-bound ligand complex structure compatible with protein-ligand (e.g., non-nucleic acid) benchmarking.
  • Composition The dataset consists of 102 (86) fragment ligands contained within 19 (15) separate (publicly available) protein complexes, of which 6 (2) and 13 (2) of these complexes are single and multi-ligand complexes, respectively. The accuracy of the dataset’s AlphaFold 3-predicted structures is measured in terms of their RMSD and TM-score (Zhang & Skolnick, 2004) compared to the corresponding ground-truth (i.e., experimental) protein structures and is visualized in Figure 10. Notably, 42% of the predicted structures have an RMSD below 4 Å and a TM-score above 0.7, indicating a portion of the dataset’s proteins have a reasonably accurate predicted structure. Given the much larger structural ensembles of this dataset’s protein complexes compared to those of the other three benchmark datasets, we believe the accuracy of these predictions may be improved with advancements in machine learning modeling of biomolecular assemblies.
  • Hosting Our preprocessed version of (the publicly available version of) this dataset (https://doi.org/10.5281/zenodo.13858866) can be downloaded from the benchmark’s Zenodo data record at https://zenodo.org/records/13858866/files/casp15_set.tar.gz.
  • Licensing We have released our preprocessed version of the (public) dataset under a CC-BY 4.0 license. The original (public) dataset is free for download via the RCSB PDB (Bank, 1971). The pre-holo-aligned protein structures predicted by AlphaFold 3 for this dataset (available on Zenodo (Morehead et al., 2024)) must only be used in accordance with the Terms of Service provided by the AlphaFold Server.
@@
Figure 11: PDBBind 2020, Astex Diverse, PoseBusters Benchmark, DockGen, and CASP15 dataset comparative analysis of protein-ligand (pocket-level) interactions.
Figure 12: Comparative analysis of the protein-ligand (pocket-level) interactions within the CASP15 dataset and baseline method predictions.
@@ F.1 Dataset protein-ligand interaction distributions
Inspired by a similar analysis presented in the PoseCheck benchmark (Harris et al., 2023), in this section, we study the frequency of different types of protein-ligand (pocket-level) interactions such as Van der Waals contacts and hydrophobic interactions occurring natively within (n.b., a size-1000 random subset of) the commonly-used PDBBind 2020 docking training set (i.e., PDBBind 2020 (1000)) as well as the Astex Diverse, PoseBusters Benchmark, DockGen, and CASP15 benchmark datasets, respectively. In particular, these measures allow us to better understand the diversity of interactions each baseline method within the PoseBench benchmark is tasked to model, within the context of each test dataset. Furthermore, these measures directly indicate which benchmark datasets are most dissimilar from commonly used training data for docking methods. Figure 11 displays the results of this analysis.
Overall, we find that the PDBBind 2020, Astex Diverse, and PoseBusters Benchmark datasets contain similar types and frequencies of interactions, with the PoseBusters Benchmark dataset containing slightly more hydrogen bond acceptors ($\sim$3 vs 1) and fewer Van der Waals contacts ($\sim$5 vs 8) on average compared to the PDBBind 2020 dataset. However, we observe a more notable difference in interaction types and frequencies between the DockGen and CASP15 datasets and the three other datasets. Specifically, we find these two benchmark datasets contain a notably different quantity of hydrogen bond acceptors and donors (n.b., $\sim$40 for CASP15), Van der Waals contact ($\sim$200 for CASP15), and hydrophobic interactions ($\sim$2 for DockGen) on average. As we will see in the DockGen benchmarking results reported in Appendix H.1, this latter observation supports our first key insight of this work, that training new docking methods on structure-based dataset clusters is a promising direction for future work on developing new pocket-generalizing docking methods.
@@
Also particularly interesting to note is the CASP15 dataset’s bimodal distribution of hydrophobic interactions, suggesting that the dataset contains two primary classes of interacting ligands giving rise to hydrophobic interactions. One possible explanation for this phenomenon is that the CASP targets, in contrast to the PDBBind, Astex Diverse, PoseBusters Benchmark, and DockGen targets, consist of a variety of both organic (e.g., drug-like molecules) and inorganic (e.g., metal) cofactors.
F.2 Protein-ligand interaction distributions of each baseline method
Intrigued by the dataset interaction patterns in Figure 11, we further investigated the (predicted) protein-ligand interactions produced by each baseline method for the (multi-ligand) CASP15 dataset, to study which machine learning-based docking method can most faithfully reproduce the true distribution of protein-ligand interactions within this benchmark dataset. Our results in Figure 12 suggest, similar to our docking results in Figure 4, that NeuralPLexer demonstrates the best overall ability to recapitulate the complex interaction dynamics observed within this dataset, presenting the unique ability (among all baseline DL methods) to correctly capture the dataset’s intricate (bimodal, top first-bottom second) interaction patterns within its hydrophobic interactions (Bogunia & Makowski, 2020; Sayyah et al., 2024). Combined with the CASP15 benchmarking results presented in Section 5 of the main text, this latter finding further supports our second key insight of this work, that the physics-informed inter-molecular clash penalties that DL methods such as NeuralPLexer employ have equipped them with physically relevant knowledge for multi-ligand docking.
Appendix G Additional method descriptions
To better contextualize the benchmark’s results comparing DL docking methods to conventional docking algorithms, in this section, we provide further details regarding how certain traditional docking methods in the benchmark leverage different sources of biomolecular data to predict protein-ligand interactions for given protein targets.
G.1 TULIP
TULIP is a template-based modeling pipeline for predicting protein-ligand interactions that we present in the benchmark as a historical reference point to better contextualize the advances of the latest DL methods for docking, as in the recent CASP15 competition template-based methods outperformed the DL docking methods that were available at the time (Xu et al., 2023). TULIP takes the target ligand’s 3D initial conformer structure (Landrum et al., 2013), the predicted receptor protein structure, and identified template structures from MULTICOM (Liu et al., 2023) as inputs. TULIP first aligns the template structures containing ligands into the same geometric space as the predicted receptor structure using UCSF Chimera’s matchmaker (Pettersen et al., 2004) in non-interactive mode. It then saves the superimposed template structures and their ligands relative to the predicted receptor structure in an output PDB file that is processed by PyRosetta’s is_ligand function (Chaudhury et al., 2010) to identify template ligands by checking each residue against the Chemical Component Dictionary of the Protein Data Bank (PDB) (Westbrook et al., 2015). The extracted unique ligands from each template and the target ligand are converted into Morgan fingerprints (Zhou & Skolnick, 2024) to compute their Tanimoto molecular similarity (Bajusz et al., 2015) (n.b., a [0, 1] metric of increasing similarity). This step provides the initial binding location of the target ligand with respect to the receptor protein structure. Furthermore, to adjust the target ligand’s binding pose and orientation by rotation and translation, TULIP uses LS-align (Hu et al., 2018) to align the target ligand with the template ligands of higher similarity through both flexible and rigid-body alignments. Between the flexible and rigid-body alignment outputs, TULIP selects the alignment with the lowest RMSD between the template and target ligands to obtain the predicted coordinates of the target ligand. Ligands with a distance greater than 6 Å from the protein surface are discarded. To handle multiple ligands with the same SMILES string, the identified ligands are grouped into n clusters, where n is the number of ligands with the same SMILES string. To compute the clusters, pairwise distances between the ligands are generated, and agglomerative clustering is used.
Appendix H Additional results
@@
H.1 DockGen results
DockGen dataset. The DockGen dataset (Corso et al., 2024a) contains 189 diverse single-ligand protein complexes, each representing a novel type of protein-ligand binding pocket. This dataset can be considered the most difficult single-ligand benchmark set since its protein binding sites are distinctly different from those commonly found in the training datasets of most deep learning-based docking methods to date.
@@ H.1 DockGen results
Benchmark results. Figures 13 and 14 reveal that DiffDock-L, RoseTTAFold-AA, and NeuralPLexer provide the best pocket generalization capabilities compared to all other baseline methods. Moreover, similar to the PoseBusters Benchmark dataset results in Section 5 of the main text, the results for DiffDock-L-Vina and P2Rank-Vina here further suggest that DiffDock-L predicts novel binding pocket locations slightly more accurately than P2Rank for conventional docking with AutoDock-Vina. Paired with the observation that ablating structural cluster training (SCT) from DiffDock yields considerably degraded DockGen performance, these findings support the idea that SCT provides DL docking methods with useful knowledge for generalizing to novel binding pockets.
Unintuitively, DiffDock-L’s results with protein-flexible relaxation applied post-prediction (i.e., DiffDock-L-Relax-Prot) demonstrate that fixed-protein relaxation (albeit unideal from a theoretical e.g., protein side chain perspective (Wankowicz et al., 2022)) yields less accuracy degradation to DiffDock-L’s original ligand conformations compared to protein-flexible relaxation. Lastly, we note that none of the baseline methods could generate any PB-valid ligand conformations, suggesting that all of their ”correct” poses are approximately accurate yet physically implausible in certain measurable ways.
@@
H.2 Expanded Astex & PoseBusters results
H.2.1 Pocket-only PoseBusters results
Figures 15 and 16 illustrate the impact of reducing the binding pocket search space of each baseline docking method by providing each method with alternative versions of the predicted PoseBusters Benchmark protein structures that have been cropped to contain only ligand-interacting ($<$ 4 Å heavy atom distance) protein residues and their (7) sequence-adjacent neighbors. Overall, we find that performing such pocket-level docking increases the docking success rates and favorably narrows the ligand RMSD distributions of DiffDock-L, DynamicBind, RoseTTAFold-AA, AutoDock Vina (w/ either DiffDock-L or P2Rank’s predicted binding pockets), and Ensemble (Con), whereas for all other baselines, performance is either maintained or degraded marginally. This finding highlights that methods such as DiffDock-L and RoseTTAFold-AA are better at leveraging a reduced (e.g., structural) search space for each ligand conformation compared to other baseline methods such as Chai-1 and NeuralPLexer.
H.2.2 Astex & PoseBusters RMSD results